OpenStudy (anonymous):

PLEASE HELP ITS CONFUSING!!!!! Simplify the expression. square root of 10-4 square root of 10

jimthompson5910 (jim_thompson5910):

hint: let \(\large x = \sqrt{10}\)

jimthompson5910 (jim_thompson5910):

so that means \[\large \sqrt{10} - 4\sqrt{10}\] turns into \[\large x - 4x\]

jimthompson5910 (jim_thompson5910):

do you see where I'm going with this?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

what does x - 4x simplify to

OpenStudy (anonymous):

3x

jimthompson5910 (jim_thompson5910):

close, but not quite

OpenStudy (anonymous):

im not sure

jimthompson5910 (jim_thompson5910):

x - 4x is really the same as 1x - 4x what is 1 - 4 equal to?

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

it's not positive though

jimthompson5910 (jim_thompson5910):

1 - 4 is -3

jimthompson5910 (jim_thompson5910):

you're subtracting in the wrong order

jimthompson5910 (jim_thompson5910):

that means 1x - 4x = -3x so x - 4x = -3x

OpenStudy (anonymous):

oh ok i get that part now

jimthompson5910 (jim_thompson5910):

and because we let \(\large x = \sqrt{10}\), we know that x - 4x = -3x turns into \[\large \sqrt{10} - 4\sqrt{10} = -3\sqrt{10}\] after you replace all the copies of 'x' with \(\large \sqrt{10}\)

jimthompson5910 (jim_thompson5910):

So here's the complete step by step picture of what's going on Let \(\large x = \sqrt{10}\) Replace each copy of square root of 10 with x. Simplify. Then once everything is simplified, replace x with the square root of 10 again. \[\large \sqrt{10} - 4\sqrt{10}\] \[\large x - 4x\] \[\large -3x\] \[\large -3\sqrt{10}\] ================================================ So, \[\large \sqrt{10} - 4\sqrt{10}\] simplifies to \[\large -3\sqrt{10}\] In other words, \[\large \sqrt{10} - 4\sqrt{10} = -3\sqrt{10}\]

OpenStudy (anonymous):

wow thanks a bunch it makes a lot more sense now!

jimthompson5910 (jim_thompson5910):

I'm glad it does